OpenStudy (anonymous):

Which conjunction or disjunction equivalent to the open sentence 4 –3|n + 2| ≥ 1?

OpenStudy (tkhunny):

Why don't you first solve the equation and then compare your correct solution to the choices? \(4 - 3|n+2| \ge 1\) \(- 3|n+2| \ge -3\) \( 3 \ge 3|n+2|\) \( 1 \ge |n+2|\) Look at each step. EXACTLY what happened? You're almost done!

OpenStudy (anonymous):

A. n + 2 ≥ 1 and n + 2 ≤ –1 B. n + 2 ≤ 1 and n + 2 ≥ –1 C. n + 2 ≤ 1 and n − 2 ≥ −1 D. n + 2 ≥ 1 or n – 2 ≤ 1

OpenStudy (tkhunny):

Perfect. Now, complete the solution and compare.

OpenStudy (anonymous):

im confused

OpenStudy (tkhunny):

It's time to get unconfused. Did you examine each step I provided? That's the first place to abandon your confusion. Tell me what happened!

OpenStudy (anonymous):

the answer is C, but I don't understand how, can you plz explain?

OpenStudy (tkhunny):

You're just not going to do it, are you? What happened in my first step? What operation was performed?